The demand for ice cream cones is P=1600 and Qd is 2 The supply of ice cream cones is P =400 and Qs is 1. The price of a cone is expressed in cents, and the quantities are expressed in cones per day. To find the equilibrium price (P*) and the equilibrium quantity (Q*), substitute Q* for QD and QS and P* for P.
1
Expert's answer
2019-04-04T09:26:52-0400

For QD, P = a - bQd and for QS, P = c + dQs, so:

1600 = a - 2b, 400 = c + d,

In equilibrium a - bQ = c + dQ,

Q = (a - c)/(b + d).